# Morphology and stellar populations of a candidate ultra-diffuse galaxy in early Euclid and Rubin imaging

### v1

@@ 1 Introduction
As imaging surveys probe larger areas of the sky at fainter flux levels, populations of LSB galaxies are being increasingly unveiled. Of particular current interest are ultra-diffuse galaxies (UDGs), with effective radii $R_{\mathrm{e}}\gtrsim$ 1.5 kpc and mean SB $\langle\mu_{g}\rangle\gtrsim 25$ mag arcsec-2 (P. G. van Dokkum et al., 2015). The formation histories of UDGs are still uncertain, and they display dramatic variations in their dark matter content – from very low to very high, with the latter hosting unusually abundant GC populations (D. A. Forbes & J. Gannon, 2024; M. L. Buzzo et al., 2025). The largest inventory of candidate UDGs is SMUDGes (Systematically Measuring Ultra-Diffuse Galaxies; D. Zaritsky et al. 2023), based on the DESI Legacy Imaging surveys (A. Dey et al., 2019). Here we present SMDG0333094--280938 at (RA, Decl = 53.2891, --28.1606, J2000), which is one of only two SMUDGes in the overlapping footprints of the first data releases from Euclid and Rubin (the other is SMDG0330098--282956). Previous work has been carried out with Euclid on cluster UDGs (T. Saifollahi et al., 2025; F. R. Marleau et al., 2025), but not yet on field UDGs, and there is so far no publication on any galaxy using Rubin data.
@@ 2 Imaging data and morphology
The Euclid imaging for SMDG0333094--280938 is from the Euclid Deep Field Fornax region of the first Quick Data Release (Q1; Euclid Collaboration et al. 2025). Figure 1 shows the Euclid image, where the galaxy has a somewhat irregular shape, including a southwest overdensity suggesting a region of recent star formation. Some point sources are visible with magnitudes of $I_{\rm E}\sim$ 25.0–26.5 and colors of $I_{\rm E}-H_{\rm E}\sim$ 0–1.0, consistent with old GCs at a distance of $\sim$ 50–60 Mpc (L. K. Hunt et al., 2025). The Legacy-based catalog size of $R_{\rm e}=6\farcs 2$ would correspond to $\sim$ 1.5–1.8 kpc at this distance, and the SB of $\langle\mu_{g}\rangle=24.6$ mag arcsec-2 would qualify this galaxy as a UDG if it were to quench and fade.
Figure 1 also shows Rubin LSSTComCam imaging from the Data Preview 1 (DP1) release in ECDFS, which has the deepest data in DP1 (NSF-DOE Vera C. Rubin Observatory, 2025), similar to the full 10-year depth of the Legacy Survey of Space and Time, with an estimated SB limit of $\mu_{g}\sim 30.3$ mag arcsec-2 and seeing of $\sim 1.1^{\prime\prime}$. The image was downloaded from the Rubin Science Platform and colorized using DS9, where the galaxy’s color and detailed morphology are readily characterized, in contrast to the Sloan Digital Sky Survey where UDGs are by definition undetected. The galaxy appears blue ($g-i\sim 0.6$) and clumpy, and some of the point sources from Euclid have GC-like optical colors ($g-i\sim 0.85$), although these could be intrinsically bluer: young stellar clusters reddened by dust (since it appears unusual for star-forming UDGs to host GCs; M. G. Jones et al. 2023).
@@ 3 Stellar population analysis and discussion
This work has made use of the Euclid Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. This publication is based in part on proprietary Rubin Observatory Legacy Survey of Space and Time (LSST) data, and was prepared in accordance with the Rubin Observatory data rights and access policies. All authors of this publication meet the requirements for co-authorship of proprietary LSST data. We thank Rubin staff for helping navigate this spectacular new dataset. This work was supported by NSF grant AST-2308390, by NASA ROSES EGIP24-0037, and by the Division of Research and Innovation at San José State University (SJSU) under Award Number 25-RSG-08-135. The content is solely the responsibility of the author(s) and does not necessarily represent the official views of SJSU.

### v2

@@ Abstract
facilities: GALEX, Rubin:Simonyi(LSSTComCam, SLAC National Accelerator Laboratory & NSF-DOE Vera C. Rubin Observatory (2024)), Euclid, Spitzer
@@ 1 Introduction
As imaging surveys probe larger areas of the sky at fainter flux levels, populations of LSB galaxies are being increasingly unveiled. Of particular current interest are ultra-diffuse galaxies (UDGs), with effective radii $R_{\mathrm{e}}\gtrsim$ 1.5 kpc and mean SB $\langle\mu_{g}\rangle\gtrsim 25$ mag arcsec-2 (P. G. van Dokkum et al., 2015). The formation histories of UDGs are still uncertain, and they display dramatic variations in their dark matter content – from very low to very high, with the latter hosting unusually abundant GC populations (D. A. Forbes & J. Gannon, 2024; M. L. Buzzo et al., 2025). The largest inventory of candidate UDGs is SMUDGes (Systematically Measuring Ultra-Diffuse Galaxies; D. Zaritsky et al. 2023), based on the DESI Legacy Imaging surveys (A. Dey et al., 2019). Here we present SMDG0333094--280938 at (RA, Decl = 53.2891, --28.1606, J2000), which is the only SMUDGe fully covered in the overlapping footprints of the first data releases from Euclid and Rubin. Previous work has been carried out with Euclid on cluster UDGs (T. Saifollahi et al., 2025; F. R. Marleau et al., 2025), but not yet on field UDGs, and there is so far no publication on any galaxy using Rubin data.
@@ 2 Imaging data and morphology
The Euclid imaging for SMDG0333094--280938 is from the Euclid Deep Field Fornax region of the first Quick Data Release (Q1; Euclid Collaboration et al. 2025). Figure 1 shows the Euclid image, colorized using DS9 (weighting each band equally). The galaxy has a somewhat irregular shape, including a southwest overdensity suggesting a region of recent star formation. Some point sources are visible with magnitudes of $I_{\rm E}\sim$ 25.0–26.5 and colors of $I_{\rm E}-H_{\rm E}\sim$ 0–1.0, consistent with old GCs at a distance of $\sim$ 50–60 Mpc (L. K. Hunt et al., 2025). The Legacy-based catalog size of $R_{\rm e}=6\farcs 2$ would correspond to $\sim$ 1.5–1.8 kpc at this distance, and the SB of $\langle\mu_{g}\rangle=24.6$ mag arcsec-2 would qualify this galaxy as a UDG if it were to quench and fade.
Figure 1 also shows Rubin LSSTComCam imaging from the Data Products 1 (DP1) release in ECDFS, which has the deepest data in DP1 (NSF-DOE Vera C. Rubin Observatory, 2025a), similar to the full 10-year depth of the Legacy Survey of Space and Time, with an estimated SB limit of $\mu_{g}\sim 30.3$ mag arcsec-2 and seeing of $\sim 1.1^{\prime\prime}$. The image was downloaded from the Rubin Science Platform and colorized using the $g,r,i$ exposures. The galaxy’s color and detailed morphology are readily characterized, in contrast to the Sloan Digital Sky Survey where UDGs are by definition undetected. The galaxy appears blue ($g-i\sim 0.6$) and clumpy, and some of the point sources from Euclid have GC-like optical colors ($g-i\sim 0.85$), although these could be intrinsically bluer: young stellar clusters reddened by dust (since it appears unusual for star-forming UDGs to host GCs; M. G. Jones et al. 2023).
@@ 3 Stellar population analysis and discussion
This work has made use of the Euclid Q1 data from the Euclid mission of the European Space Agency (ESA), 2025 hosted at IPAC (Euclid Consortium, 2025). This publication is based in part on proprietary Rubin Observatory Legacy Survey of Space and Time (LSST) data (NSF-DOE Vera C. Rubin Observatory, 2025b), and was prepared in accordance with the Rubin Observatory data rights and access policies. All authors of this publication meet the requirements for co-authorship of proprietary LSST data. We thank Rubin staff for helping navigate this spectacular new dataset. Other datasets used include the Bianchi, Conti, Shiao (BCS) Catalog of Unique GALEX Sources (L. Bianchi et al., 2014) hosted at MAST (L. Bianchi, 2014), and the Spitzer Level 2 / post Basic Calibrated Data hosted by IPAC (Spitzer Science Center, 2020). This work was supported by NSF grant AST-2308390, by NASA ROSES EGIP24-0037, and by the Division of Research and Innovation at San José State University (SJSU) under Award Number 25-RSG-08-135. The content is solely the responsibility of the author(s) and does not necessarily represent the official views of SJSU.